# An LLM-based Agent for Reliable Docker Environment Configuration

### v2

@@ 3.1 External environment
$\bullet$ Action-observation interaction: In each turn of the environment-agent interaction, the event history $E_{t}={(a_{1},o_{1}),(a_{2},o_{2}),\ldots,(a_{t-1},o_{t-1})}$ records the historical actions and observations up to time step $t$. Additionally, the sequence of successfully executed commands ${C_{1},C_{2},\ldots,C_{n}}$ helps remind the LLM agent of the current system state.
At time step $t$, given the event history $E_{t}$ and the sequence of successfully executed commands $C_{1},C_{2},\dots,C_{n}$, the LLM agent generates a response including an action $a_{t}$:
$a_{t}=\pi(E_{t},{C_{1},C_{2},\ldots,C_{n}})$ (4)
where $\pi$ is the function of the LLM agent’s decision. The environment state then transitions based on the action $a_{t}$:
$S_{t+1}=\delta(S_{t},a_{t})$ (5)
This iterative process ensures that the LLM agent remains informed of the system’s current state and makes decisions accordingly, facilitating effective multi-turn interaction within the environment.
$\bullet$ Rollback ($\rho$): As illustrated in Figure 1, commands $C_{fail}$ may lead to an uncertain state ($S_{u}$) if not successfully executed. When a command executes failed (i.e., the command returns a non-zero return code), the system transitions into an uncertain state as defined by:
$S_{u}=\delta(S_{k-1},C_{fail})$ (6)
Here, $C_{fail}$ is the failed executed command. If a command fails, the system will transition to the uncertain state $S_{u}$. To avoid such uncertainty, we introduce rollback ($\rho$). In the event of a failure at any step, the rollback mechanism restores the environment to the last certain state:
$\rho:\mathcal{S}\times\mathcal{C}\rightarrow\mathcal{S},\quad\rho(S_{u},{C_{1},C_{2},\ldots,C_{k}})=S_{k-1}$ (7)
Specifically, each time an instruction ($C$) is executed, we utilize the “docker commit” command to take a snapshot of the current state. Once the instruction is completed, we check its return code. If the return code is not 0, it indicates a failure in executing the instruction, and we replace the image with the most recently committed one. Besides, some instructions, such as “cat”, generally do not change the state of the environment, and thus, we do not apply the rollback mechanism to them. The complete list of these actions is shown in Appendix B.
@@
$\bullet$ Base image change: If the LLM agent identifies that the currently selected base image is incorrect or unsuitable during the environment configuration process, it can reselect the base image. Once a new image is chosen, all previous configurations become invalid, and the executed commands are cleared, necessitating a restart of the configuration process. For instance, as shown in Figure 3, when the LLM agent discovers that the “StrEnum” class cannot be imported through testing, which is only available in the Python standard library starting from version 3.11, it decides to reselect the base image to python:3.11.
$\bullet$ Result processor: In each turn of the interaction, executing commands may generate extensive output (e.g., logs), which can greatly increase the burden of the LLM agent. Therefore, whenever such information is added to the event history, we truncate excessively long text, retaining only the initial and final sections up to a specific length.
@@
$\bullet$ Dockerfile generator: If the LLM agent successfully runs all tests in the repository in the internal environment, the configuration process needs to be recorded in a Dockerfile. Figure 4 illustrates the Dockerfile generator’s operation. 4 (a) lists the internal commands actually executed in the internal environment, their return codes (e.g., 0 indicates success, otherwise failure), and the current directory (dir). 4 (b) shows the process of generating the Dockerfile.
@@
The principles for the conversion process include the following steps, scanning all internal commands sequentially, and the line numbers mentioned below match the line number of Figure 4 (a):
@@
    For commands that run successfully (i.e., have a return code of 0), prepend the command with “RUN” to form a Dockerfile statement (e.g., line 3). Commands that fail (i.e., have a return code other than 0, such as line 4) are rolled back and not included in the Dockerfile. An exception is some commands typically do not change the current state (such as “cat” on line 2). These commands are not added to the Dockerfile. We show the details in Appendix B.
@@
    If a base image change statement is encountered (e.g., line 5), discard all previous configurations and switch to the new base image.

### v3

@@ 3.1 External environment
The external environment serves as a bridge between the LLM and internal environment. It transmits actions from the LLM to the internal environment and relaying observations back. Following with the ReAct framework [13], the external environment maintains a history of thoughts, actions (i.e., successfully executed commands), and observations. This iterative process ensures that the LLM stays informed about the system’s state and makes accurate decisions while effectively interacting with the internal environment.
$\bullet$ Rollback: As shown in Figure 3, commands like “commandA”  [^2] We collected some commands that can “pollute” the environment, see Appendix C. may “pollute” the environment, making it difficult to reproduce using a Dockerfile. When a command fails (i.e., returns a non-zero return code), the environment transitions to an uncertain state. Therefore, we introduce a rollback action to ensure the building process remains consistent. Specifically, before executing a command, we use “docker commit” to create a reserve snapshot of the current state (i.e., save the image at that point). If the return code is not 0, indicating a failure, the environment is rolled back to the most recent reserve snapshot. However, certain commands, referred to as safe commands (e.g., “cat”), generally do not alter the state of the environment and are exempt from rollback. For a complete list of safe commands, see Appendix B.2.
$\bullet$ Base image change: If the LLM agent determines that the current base image is unsuitable during the environment building process, it can reselect and switch to a new base image. This change invalidates the previous building process, clearing all executed commands and requiring a restart.
$\bullet$ Result processor: Command execution during interactions can produce extensive output (e.g., error logs), potentially overwhelming the LLM agent. To mitigate this, long outputs are truncated by this action, retaining only the initial and final sections up to a specified length.
3.2 Internal environment
The internal environment is a Docker container. Based on the latest data [14] from 2025, we select Python 3.10 as the default Docker base image due to its broadest adoption among Python versions.
$\bullet$ Environment monitoring: It serves as the eyes of the LLM agent in the internal environment, allowing it to observe the current environment state. These commands typically do not change the state of the environment. Basic commands like “ls” and “cat” are used to inspect directories and files, “find” is used to locate files, while more advanced commands like “pip list” and “pipdeptree” [15] retrieve the versions of installed libraries and dependency relationships.
$\bullet$ Dependency installation: It serves as installing third-party packages required for running tests, using tools like pip for Python and apt-get for system-level packages. To resolve potential conflicts (e.g., version constraints), a dependency management is implemented to assist the installation, where packages to be installed are added to a waiting list and resolved as needed. If installation succeeds, the environment is updated; otherwise, the rollback is performed. For detailed designs, see Appendix F.
$\bullet$ Test running: It serves as both a compass and a checkpoint, guiding the building process and verifying whether the environment runs all tests successfully. By executing “pytest” for unit tests, it determines whether the Docker container has been correctly built. If all tests run successfully, the process concludes; otherwise, error logs are sent to the LLM agent for further adjustments.
$\bullet$ Code editing: It enables the LLM agent to modify the code within the internal environment, including both inside and outside the repository. Direct code editing is rare, but sometimes needed for issues like syntax errors. To prevent the LLM agent from bypassing tests by directly altering or deleting test files, it is restricted from modifying or deleting the original test files within the repository.
$\bullet$ Bash commands: Like many other LLM-based agents [2, 3, 4], Repo2Run is allowed to invoke bash commands, allowing it fully operate within the Docker container. It enables the agent to execute dynamically defined actions that are not pre-specified, ensuring flexibility in handling various errors.
Figure 5: Rules for Dockerfile synthesis, illustrating how the Dockerfile synthesizer maps executed commands into Dockerfile statements using four keywords: “FROM”, “ENV”, “COPY”, and “RUN”. Black arrows represent the creation of statements, while green arrows indicate their transformations. Red text next to the arrows specifies the commands executed during each step.
3.3 Dockerfile synthesizer
Once the LLM agent successfully runs all tests, the Dockerfile synthesizer converts the command sequence from the building process into a runnable Dockerfile. It processes each command sequentially, following the rules in Figure 5 to synthesize Dockerfile statements. The synthesis uses four Dockerfile keywords: “FROM”, “ENV”, “COPY”, and “RUN”. For more details and examples, see Appendix B.
$\bullet$ FROM: The FROM statement defines the base image and is typically the first line in a Dockerfile. If the LLM agent changed the base image, the FROM statement must be updated, and all subsequents are cleared to ensure consistency.
$\bullet$ ENV: The ENV statement persistently sets environment variables in the Docker container. When a command with “export” is detected, it is converted into an ENV statement. If the vaule of an environment variable is overwritten, this statement is updated accordingly.
$\bullet$ COPY: The COPY statement copies local files or directories into the Docker container. As shown in Figure 5, it is often used for code editing to import editing scripts (i.e., code_edit.py) and patches.
$\bullet$ RUN: The RUN statement executes commands in the container, with each line creating a new bash session. It cannot be used for persistent environment variables. Specially, for every installed packages, versions are recorded to ensure reproducibility, and the RUN statement is updated accordingly.
4 Experiment
We evaluate the effectiveness of Repo2Run on 420 Python code repositories. As the popular option, we select gpt-4o-2024-05-13 for all experiments, with the temperature uniformly set to 0.2.
4.1 Benchmark
To the best of our knowledge, there is no prior work similar to Repo2Run that automates executable environment building. Existing manually constructed datasets are limited to very few repositories [1]. To validate the capability of Repo2Run, we create a new benchmark consisting of filtered Python repositories from GitHub based on the following criteria:
$\bullet$ Creation date: To avoid the data leakage, we select repositories created in 2024, ensuring they are not part of mainstream LLM training data.
$\bullet$ Star count: To maintain quality, we only include repositories with more than 100 stars.
$\bullet$ Test directory: We focus on repositories likely to contain unit tests, identified using pytest, a leading Python testing framework compatible with tools like unittest. Pytest detects test files with a “test_” prefix or “_test” suffix, typically located in “test” or “tests” directories. We only retain repositories that have these directories.
Using these criteria, we crawled 449 repositories in December 2024 and filtered 420 containing at least one unit test to form our benchmark. For statistic of their scale, see Appendix D.
4.2 Evaluation metrics
$\bullet$ Dockerfile Generation Success Rate (DGSR): It indicates the percentage of attempts where the method successfully generates a runnable Dockerfile. To be considered successful, the generated Dockerfile must be able to build without errors. If the Dockerfile for a code repository successfully builds, it is regarded as a successful generation. Generating runnable Dockerfile is fundamental for successfully building the executable environment.
$\bullet$ Environment Building Success Rate (EBSR): It represents the percentage of attempts where the method successfully builds executable environments. For a successful building, the generated Dockerfile must not only build successfully but also allow tests to run by “pytest” in the Docker container. We are only concerned with whether tests can be executed, regardless of whether they pass or fail, as outcomes of tests may inherently vary within the repository.
4.3 Baselines
$\bullet$ pipreqs [16]: It is an automated tool that generates a “requirements.txt” file by analyzing the import statements in the Python scripts and identifying the necessary dependencies without LLM. Using the requirements.txt file generated by pipreqs, we create a Dockerfile. The detail is provided in Appendix H.1.
$\bullet$ LLM generator: The “README” file in a code repository usually contains environment building instructions. Therefore, we directly drive the LLM to read the “README” file and generate an executable Dockerfile accordingly.
$\bullet$ SWE-agent [3]: SWE-agent establishes a custom agent-computer interface (ACI) that uses the LLM agent’s interaction with the repository environment by allowing actions such as reading files, editing files, and executing bash commands. Initially intended as an LLM agent for bug fixing, we preserve its framework and default settings, adjust its prompts, as shown in Appendix H.2.
4.4 Experimental Results
Table 1: The results of different baselines, including the percentage and number of successfully generated Dockerfiles and successfully built environments.
Metric DGSR # Successfully Generated Dockerfiles ECSR # Successfully Built Environments
pipreqs 29.8% 125 6.0% 25
LLM generator 47.6% 200 22.1% 93
SWE-agent 26.9% 113 9.0% 38
Repo2Run 100% 420 86.0% 361
Table 2: Repository distribution across different domains and their EBSR, where SE means Software Engineering, NLP means Natural Language Processing, and DV means Data Visualization.
Type AI / ML SE NLP DV Security Others All
# Success 221 70 41 10 7 12 361
# Total 267 73 47 11 8 14 420
EBSR 82.8% 95.9% 87.2% 90.9% 87.5% 85.7% 86.0%
The results of different baselines are presented in Table 1. We observe that Repo2Run consistently outperforms other baselines on both DGSR and EBSR. Repo2Run ultimately completed environment building for 361 code repositories, achieving an EBSR of 86.0%. It is 63.9% higher than the highest rate achieved by other methods, demonstrating great advantages. Due to the design of rollback, Repo2Run successfully generates Dockerfiles that can be built successfully for all 420 code repositories, which other actions cannot guarantee. For details of each repostories, see Appendix G.
For pipreqs, the main failures come from two reasons. First, generating the requirements.txt fails when there are issues within the repository, such as encoding errors or syntax errors in the files. This happens in 30 repositories (7.1%). Second, even when requirements.txt is generated, it might not be downloaded properly due to package version conflicts. This occurs in 265 repositories (63.1%). Besides, both the LLM generator and SWE-agent fail to ensure that the generated Dockerfile can be successfully built due to the lack of an ensuring mechanism. Surprisingly, the ability of SWE-agent, a general agent framework, to generate Dockerfiles is even weaker than simply letting the LLM read the “README” file. This indicates that a general agent framework cannot guarantee the generation of runnable Dockerfiles. Ensuring mechanisms like rollback is necessary to effectively use the interactive information from the agent to generate runnable Dockerfiles.
As shown in Table 2, we manually categorized the 420 repositories into six domains. EBSR across these domains all exceed 80%, outperforming all baselines and demonstrating Repo2Run’s consistent performance across different domains. Prior research [17, 18] shows that the real-world benchmarks differs from synthetic ones and better reflect a model’s true capabilities. Our benchmark is built entirely from real-world GitHub data, filtered only by repository creation date, star count, and test, with no additional filtering. Thus, it accurately represents the current distribution of Python repositories. Moreover, a previous real-world benchmark [18] includes 59.6% of AI/ML repositories, closely matching the 63.6% proportion in our benchmark. This alignment further confirms that our benchmark reflects the real-world scenarios.
4.5 Ablabtion of Repo2Run
Table 3: The results of ablation experiments, including the percentage and number of successfully generated Dockerfiles and successfully built environments.
Metric DGSR # Successfully Generated Dockerfiles ECSR # Successfully Built Environments
w/o dual-environment 92.4%  $\downarrow\,7.6$ 388 41.7%  $\downarrow\,44.3$ 175
$\hookrightarrow$ w/o rollback 96.9%  $\downarrow\,3.1$ 407 83.6%  $\downarrow\,2.4$ 351
w/o Dockerfile generator 19.5%  $\downarrow\,80.5$ 82 13.8%  $\downarrow\,72.2$ 58
Repo2Run 100% 420 86.0% 361
To investigate the impacts of the dual-environment architecture and Dockerfile synthesizer separately, as two parts of Repo2Run, we separately remove each component of them. For the experiment without the dual-environment architecture, we retain only the internal environment’s bash commands as the most basic interface and remove all other actions. For the experiment without the Dockerfile synthesizer, we directly instruct the LLM to synthesize a runnable Dockerfile.
Experimental result of the ablation study is shown in Table 3. We observe that removing the dual-environment architecture and retaining only bash commands results in a 7.6% decrease in DGSR. The main reason for this drop is the removal of rollback and other designs, making the system more prone to entering uncertain states and subsequently failing to reproduce. In addition, EBSR shows a 44.3% decrease, primarily because the simplification of design makes it more difficult for the LLM agent to execute all tests in the internal environment. Besides, removing the Dockerfile synthesizer directly leads to an 80.5% drop in DGSR. This indicates that having the LLM directly generate Dockerfiles is unlikely to fully follow the event history, resulting in Dockerfiles that fail to build successfully. This also directly causes a sharp decline in EBSR. Specially, we also perform an ablation study for the rollback. The results show that removing the rollback causes 3.1% of generated Dockerfiles to become unrunnable, demonstrating the effectiveness of the rollback mechanism.
It is also observed that Repo2Run without the Dockerfile synthesizer is outperformed by the LLM generator. This is because the LLM generator leverages the “README” file, which provides a clear, simple and high-level overview of the executable environment building, allowing for more accurate Dockerfile synthesizer. In contrast, the event history-based approach lacks this context, making it harder for the LLM to fully understand the building goals. However, the Dockerfile synthesizer effectively utilizes the detailed event history, highlighting the complementary roles of both components of Repo2Run in generating a reliable runnable Dockerfile.
5 Discussion
Figure 6: Time distribution of successful building.
5.1 Time consumption
Figure 6 shows the distribution of time spent successfully building 361 code repositories. The average time for successful building using Repo2Run is 29.03 minutes. 111 (30.7%) of the repositories are successfully built in less than 10 minutes. Additionally, our empirical study through sampling indicates an average manual building time of 21.33 minutes (Appendix E). Considering network differences and randomness, Repo2Run achieves a time consumption comparable to manual building. Additionally, for complex issues, Repo2Run shows greater advantages over manual building. Repo2Run successfully builds all the cases that were manually successful in our empirical study. Moreover, the cases where Repo2Run fails are also not successfully built manually.
Table 4: Analysis of failed cases in Repo2Run.
Category # Case (%)
Hardware Insufficiency 21 (35.6%)
Missing Token 1 (1.7%)
Repository Defects 5 (8.5%)
Dependency Installation Timeout 8 (13.5%)
Runtest Timeout 24 (40.7%)
Table 5: Comparison of Dockerfile quality. “Warning” counts in each group show Repo2Run vs. baseline.
Baseline # “Warning”
vs. pipreqs 5.17 / 6.0
vs. LLM generator 5.42 / 10.16
vs. SWE-agent 5.43 / 6.49
5.2 Failure case study
For the code repositories that fail to build, we manually inspect the reasons for failures and find that most are due to issues within the repositories themselves. Table 4 demonstrates the analysis of 59 failed cases in Repo2Run. The first three categories (45.8%) primarily fail due to external factors, while the latter two categories (54.2%) are unable to be configured successfully within the allotted time due to excessive dependencies and extensive tests inherent to the repositories. We also provide a detailed case in Appendix I.
5.3 Dockerfile Quality
We utilize hadolint [19] to analyze the quality of the Dockerfile generated by Repo2Run, comparing with Dockerfiles successfully generated by three baselines and counting the number of warnings. Fewer warnings indicate higher quality in Dockerfile generation. As shown in Table 5, Dockerfiles generated by Repo2Run’s consistently have fewer warnings than those generated by baselines. Specifically, compared to pipreqs, LLM generator, and SWE-agent, Repo2Run reduces warnings from 6.0 to 5.17, 10.16 to 5.42, and 6.49 to 5.43, respectively. The results indicate that Repo2Run  produces higher-quality Dockerfiles with fewer potential issues.
5.4 Limitation
Repo2Run can automate building executable environments for many repositories but may still encounter challenges in some edge cases. For complex environments that remain unresolved, manual verification and intervention are necessary. Nonetheless, as an important step toward full automation, we believe this work makes a significant contribution to the AI community’s infrastructure.
6 Related Work
6.1 Environment building and Dockerfile generation
Oss-Fuzz-Gen [20] relies on predefined build instructions (such as “./bootstrap.sh”, “./configure”, “make”) to build projects for fuzzing but lacks of flexibility for diverse projects when specified files are absent. Existing solutions that help developers write Dockerfiles broadly fall into three categories: (1) Template-based generators that create Dockerfiles based on project context [21, 22], (2) Task-specific tools such as pipreqs [16], DockerizeMe [23] which supports environment dependency inference for Python projects and DockerGen  [24] for dependency recommendations based on knowledge graphs built from existing Dockerfiles, and (3) Code completion tools including GitHub Copilot [5] and HumpBack [25] that generate suggestions for developers while writing Dockerfiles. The use of deep learning models to generate Dockerfiles based on natural language specifications of software requirements is also investigated [26]. While these approaches provide valuable assistance, they either require significant manual input from developers or are limited to specific use cases.
6.2 LLM-based agent
LLM-based agents typically consist of four key components: planning, memory, perception, and action [27]. Planning is crucial for agent systems, as it schedules agents to ensure a smooth process. LLM-based agents employ various planning strategies, including single [28] or multiple planners [29], single [30] or multi-turn planning [31], and single [32] or multi-path planning [33]. The memory component in LLM-based agents stores historical data to support coherent reasoning and complex tasks. Implementations vary in terms of memory duration (short-term [34] or long-term memory [35]), memory ownership (specific [36] or shared memory [37]). For perception, LLM-based agents primarily utilize textual input [38, 35] (natural language and programming language) and visual input [39, 40] (images and diagrams) to perceive and process information. To extend capabilities beyond interactive dialogue, the action component employs various external tools [41, 42], such as searching tools, file operations, and GUI operations. Nowadays, LLM-based agents have demonstrated superior performance compared to standalone LLMs in various software engineering tasks [43, 44, 45, 46, 47]. However, no LLM-based agents are specifically designed for executable environment building currently. To fill this gap, this paper employs a novel approach for automated executable coding environment building and runnable Dockerfile generation.
7 Conclusion
In this paper, we propose Repo2Run, the first LLM-based agent for automated coding executable environment building and Dockerfile generation for Python repositories. With a dual-environment architecture and a Dockerfile synthesizer, Repo2Run is able to select and change the base image, manage and install dependencies based on action observation and rollback mechanism, and utilize bash commands and existing test suites. Our evaluation of 420 popular Python repositories hosted on GitHub demonstrates the effectiveness of Repo2Run with an 86.0% success rate. We believe Repo2Run will serve as the foundational infrastructure, enabling the community to efficiently scale up executable environments.
Appendix
Appendix A Repo2Run details
A.1 Example of Repo2Run
Figure 7 shows an example process of Repo2Run, including the event stream and Repo2Run  workflow. Below, we introduce the executed action based on the example:
@@ A.1 Example of Repo2Run
  • Base Image Change: In the tenth part of Event Stream, when the LLM agent discovers that the “StrEnum” class cannot be imported through testing, which is only available in the Python standard library starting from version 3.11, it decides to reselect the base image to python:3.11.
  • Environment monitoring: In the fourth part, the LLM agent calls the “find” command to search for requirements.txt within the environment to determine which dependencies to install.
  • Dependency installation: In the sixth part, the action “waitinglist addfile” adds all elements from requirements.txt to the waiting list, while executing the “download” command. Third-party libraries are taken one by one from the waiting list and installed in the internal environment. The installation commands (e.g., “pip install” and “apt-get install”) are executed.
  • Code editing: The eighteenth part demonstrates code editing to correct a syntax error resulting from the improper use of double quotes in a Python f-string.
  • Bash commands: The sixteenth part shows that the LLM agent uses the command “export PYTHONPATH=/repo/src”.
  • Test running: The eighth part shows that the LLM agent calls the “runtest” command to execute unit tests.
Figure 7: An example process of Repo2Run, which illustrates two main parts: Event Stream and Repo2Run Workflow. The Event Stream tracks the entire action-observation process, where the green box represents the abbreviated system prompt, yellow boxes represent responses from the LLM agent including the actions, and blue boxes represent observations from the Runtime Environment. The shaded actions indicate buildings abandoned after base image change. In the Event Stream, the blue text indicates that the command starts running, the green text indicates that the command runs successfully, and the red text indicates that the command fails. The Repo2Run Workflow consists of the internal environment and the external environment. The internal environment serves as the actual buildings Docker-based sandbox, which builds an actual testing runtime environment. The external environment executes the action-observation process and assists in the building process within the internal environment.
A.2 Action usage frequency
Figure 8: Action usage frequency of the success building.
As shown in Figure 8, we analyze the invocation times of various action types in 361 successfully built projects, including the five actions within the internal environment and the action of base image change. Bash commands are the most frequently used action, as they encompass the majority of instructions. Additionally, we observe that the LLM agent tends to call dependency installation quite frequently, averaging about 18 times per building, which means roughly 18 dependencies are installed per building on average. Moreover, the LLM agent calls test running approximately 3.5 times per building on average, which typically helps the agent better identify issues. Among successful buildings, we find 48 instances of changing the base image, accounting for 13.3%  of success cases, indicating that the initial selection of the base image is often incorrect and requires subsequent adjustments.
Appendix B Dockerfile synthesizer details
B.1 Example
Figure 9: An example of Dockerfile synthesizer to transfer the commands into a runnable Dockerfile.
@@ B.1 Example
If the LLM agent successfully runs all tests in the repository in the internal environment, the building process needs to be recorded in a Dockerfile. Figure 9 illustrates the Dockerfile synthesizer’s operation. Figure 9 (a) lists the internal commands actually executed in the internal environment, their return codes (e.g., 0 indicates success, otherwise failure), and the current directory (dir). Figure 9 (b) shows the process of generating the Dockerfile.
@@
The principles for the conversion process include the following essential steps: scanning all internal commands sequentially, and the line numbers mentioned below precisely match the line number of Figure 9 (a):
@@
    For commands that run successfully (i.e., have a return code of 0), prepend the command with “RUN” to form a Dockerfile statement (e.g., line 3). Commands that fail (i.e., have a return code other than 0, such as line 4) are rolled back and not included in the Dockerfile. An exception is that some commands typically do not change the current state (such as “cat” on line 2). They are not added to the Dockerfile. We show the complete list in Appendix B.2.
@@
    If a base image change statement is encountered (e.g., line 5), discard all previous buildings and switch to the new base image.